Whitman has a direct labor standard of 2 hours per unit of output. Each employee has a standard wage rate of $28.50 per hour. During July, Whitman paid $190,200 to employees for 8,920 hours worked. 4,710 units were produced during July. What is the direct labor efficiency variance?

$14,250 favorable

$64,020 favorable

$64,020 unfavorable

$78,270 favorable
Answer #1

Labor efficiency variance = (SH-AH) *SR

= (4710*2 - 8920) * 28.50

= (9420 - 8920) * 28.50

= 14,250 Favourable

Option A is the answer